OpenStudy (anonymous):

the graph of y = x^3 -9x^2 - 16x +1 has a slope of 5 at two points. Find the coordinates of the points

OpenStudy (asnaseer):

Find the two values of x that satisfy dy/dx = 5. Then substitute these x values into your original equation to find the corresponding y values.

OpenStudy (anonymous):

Differentiate y = x^3 -9x^2 - 16x +1 dy/dx = 3 x^2 -18x-16 Slope of 5. Therefore, dy/dx = 5 5= 3 x^2 -18x-16 3 x^2 -18x-21 =0 Solving for x, by factorising, (3x-21)(x+1) = 0 Values of x= 7 and -1 To find the 2 points, solve for y by substituting x = 7 and -1 For x = 7 y = 7^3 -9(7)^2 - 16(7) +1 y=-209 (7,-209) For x = -1 y = (-1)^3 -9(-1)^2 - 16(-1) +1 y=7 (-1,7)